ASTROCHEMISTRY WITH THE UPGRADED COMBINED ARRAY FOR RESEARCH IN MILLIMETER-WAVE ASTRONOMY

\underline{D. N. FRIEDEL}, Department of Astronomy, University of Illinois, Urbana, IL 61801.

The Combined Array for Research in Millimeter-wave Astronomy (CARMA) has been actively observing at λ=1mm, 3mm, and 1cm frequencies (210-270 GHz, 85-115 GHz, and 30 GHz) for several years. This upcoming year will see many new improvements to the system, including a new correlator and dual polarization receivers. This talk will introduce the new, upgraded imaging capabilities of this state of the art instrument.